# Charge-ordered states in twisted MoTe2

### 05 Mar 2026 v1

@@ Abstract
The recent observations of fractional quantum anomalous Hall effects (FQAHEs) in twisted MoTe2 (tMoTe2) homobilayers [6, 40, 55, 59, 22] and in multilayer rhombohedral graphene [29, 36, 28] aligned with hexagonal boron nitride, have established these systems as platforms for fractional Chern insulator (FCI) states. FCIs emerge when strong electron–electron interactions dominate the kinetic energy within a topologically nontrivial, nearly flat electronic band [35, 45, 52, 44, 20, 4, 19]. Like their magnetic-field-driven counterparts — fractional quantum Hall effect (FQHE) [49] states — FCIs compete closely with interaction-driven phases that break translational [42, 24]. In the FQHE case, electron and hole Wigner crystals [53], bubble phases [16], and stripe phases [27] are well-established as ground states over some ranges of Landau-level filling [57, 58, 31, 9, 5, 30, 10]. The recent observation of reentrant integer quantum Hall states [56, 29, 38] in both rhombohedral graphene and TMD homobilayer systems suggests that broken translational symmetry states are also ground states at some fractional band filling factors in these systems.
@@
Figure 1: Flat bands, real-space structure and effective moiré potential of tMoTe2. (a) Bandwidth (meV) versus twist angle of the Adiabatic model (first-shell approximation) and the Continuum model for tMoTe2. (b) Schematic of the moiré unit cell illustrating MX, XM and MM sites, and $\sqrt{3}\times\sqrt{3}$ and $2\times 2$ commensurate supercells. (c) Amplitude of the leading (first-shell) harmonic of the effective moiré potential $V_{1}(\theta)$ versus twist angle; $V_{1}$ changes sign at $\theta_{c}\simeq 3.7^{\circ}$, switching the potential minima from MX (or XM, respectively) on the unit cell edges to $M$ at the unit-cell center. (d) The Landau-level mixing parameter, $\kappa\equiv[e^{2}/\epsilon\,\ell_{B}]/(\hbar\omega_{c})$, versus twist angle for $\varepsilon=10.0$. In (a,d), the vertical gray line marks $\theta_{c}\simeq 3.7^{\circ}$. (e) Illustration of how the valence-band holes tend to localize on the MX or XM sites for $\theta<\theta_{c}$ and on the MM sites for $\theta>\theta_{c}$ in tMoTe2. We use model parameters from Ref. [51].
@@
where $\hat{\bm{\Pi}}=-\mathrm{i}\hbar\nabla+e\bm{\mathcal{A}}_{0}/c$. The scalar potential $U_{\mathrm{ad}}(\bm{r})$ collects contributions from $|\bm{\Delta}(\bm{r})|$, $\Delta_{0}(\bm{r})$, and the quantum metric of the pseudospin texture. The uniform component $\mathcal{B}_{0}$ is conveniently treated by using the corresponding Landau level representation for $H_{\mathrm{ad}}$. The field $\mathcal{B}_{0}$ defines an effective magnetic length $\ell_{B}=\sqrt{\hbar c/(e\,\mathcal{B}_{0})}$ and cyclotron energy $\hbar\omega_{c}=e\,\mathcal{B}_{0}/(m^{*}c)\propto\theta^{2}$ . We quantify the interaction strength by the LL-mixing parameter $\kappa\equiv[e^{2}/(\varepsilon\,\ell_{B})]/(\hbar\omega_{c})\propto 1/\theta$. When projected to the lowest Landau level (LLL), $V$ acts as an effective scalar potential $V(\bm{r})$, see Supplemental Material [3] for details. We denote the amplitude of its leading (first-shell) Fourier harmonic by $V_{1}(\theta)$, see Fig. 1(c). The sign change of $V_{1}$ at the critical angle $\theta_{c}$ reverses the landscape of potential minima and therefore selects the preferred CDW pinning site.
Interactions and Hartree–Fock theory.— We study the influence of long-range Coulomb interactions, $H_{\mathrm{C}}=\frac{1}{2A}\sum_{\bm{q}}V_{q}:\hat{n}_{-\bm{q}}\hat{n}_{\bm{q}}:$, using self-consistent Hartree–Fock (HF) theory, which provides an accurate description of CDW states at fractional band fillings. We employ the Landau gauge $\bm{\mathcal{A}}_{0}=\mathcal{B}_{0}\,x\,\hat{\bm{y}}$, in which states within a Landau level are labeled by guiding center $X$. The mean-field Hamiltonian is
@@
CDW properties.— When the moiré potential is ignored, our problem reduces to that of holes in Landau levels for which the CDW states are hole Wigner crystals (one hole per unit cell) when $\nu_{h}<1/2$ and electron Wigner crystals for $\nu_{h}>1/2$  [^1] We work in a Hilbert space in which the microscopic degrees of freedom are holes. By electron Wigner crystal we mean a state in which the CDW unit cell area contains one hole less than a full Landau level. When Landau levels mixing is weak, these states are expected to be stable for $\nu_{h}>1/2$ because electrons are then more readily localized than holes.. The HF ground-state energies $E_{\mathrm{tot}}(\theta)$ we obtain for $\nu_{h}=1/3,\,2/3,\,1/4,\,1/2,$ and $3/4$ demonstrate that the moiré potential influences the CDW ground state. We first discuss $\nu_{h}=1/3$ and $2/3$, for which we consider $\sqrt{3}\!\times\!\sqrt{3}$ CDW supercells containing, respectively, one and two holes.
@@
Figure 2: Competing charge-ordered states over a range of hole fillings and twist angles. (a) Total energy per hole $E_{\textrm{tot}}$ versus twist angle $\theta$ for competing phases at fillings $\nu_{h}=1/4,\,1/3,\,1/2,\,2/3,\,3/4$. The line style indicates the rotational symmetry of the corresponding state ($C_{6}$, $C_{3}$, $C_{2}$); first-order transition angles are marked by square markers. Numbered markers label configurations whose real-space densities are shown in (b). (b) Real-space hole density $\rho_{h}(\bm{r})$ for the six labeled configurations in (a). White hexagons mark the moiré unit cell and red hexagons mark the CDW unit cell. Single particle energies in (a) are relative to the twist-angle-dependent moiré band energy at $\nu_{h}=0$.
At $\nu_{h}=2/3$ the trend reverses: two holes fill each $\sqrt{3}\!\times\!\sqrt{3}$ CDW unit cell as illustrated in Fig. 2(a). For $\theta<\theta_{c}$, holes occupy both MX and XM sites and form a $C_{6}$-symmetric state, leaving MM sites to be occupied by electrons (missing holes). (See Fig. 2(b), panel 4.) For $\theta>\theta_{c}$, the moiré potential peaks at MX and XM sites and drives hole accumulation at MM, and either an MX or an XM site, lowering the symmetry from $C_{6}\!\to\!C_{3}$, corresponding to electrons at XM or MX sites (See Fig. 2(b), panel 5.) [1]. Note that the energy difference between these crystal motifs is much smaller at $\nu_{h}=2/3$ compared to at $\nu_{h}=1/3$. We attribute this difference to a reduction in the lattice-pinning energy in the electron crystal case that cannot take advantage of Landau level mixing  [^2] At $\nu_{h}=1/3$ the crystal is a hole Wigner crystal in which the hole density $\rho_{h}(\bm{r})$ is maximized at lattice sites and increases with the Landau-level mixing parameter $\kappa$ (see Fig. 1(d)). In contrast, at $\nu_{h}=2/3$ the competing state is an electron CDW, corresponding to fully depleting $\rho_{h}(\bm{r})$ at lattice sites. Because $\rho_{h}(\bm{r})\!\geq\!0$, the hole density at lattice sites is limited and LL mixing is less effective in further lowering energies..
@@
Figure 3: (a) Lowest-Landau-level weight $w_{\mathrm{LLL}}$ of the self-consistent CDW solutions versus twist angle $\theta$ (see SM [3] for the definition of $w_{\mathrm{LLL}}$). The light-blue band shows the range across fillings $\nu_{h}\in\{1/4,1/3,1/2,2/3,3/4\}$ as a function of $\theta$. Representative curves are overlaid: color encodes filling $\nu_{h}$, and line style encodes symmetry ($C_{6}$, $C_{3}$, $C_{2}$). The light-gray curve traces the minimum $w_{\mathrm{LLL}}^{\mathrm{min}}(\theta)$ across all datasets. (b) Total energy per hole versus filling at representative twist angles $\theta=3.0^{\circ}$, $3.7^{\circ}$, and $4.2^{\circ}$; shaded bands cover the energy range over $\theta\in[2.5^{\circ},4.5^{\circ}]$. Marker style indicates when $C_{6}$ symmetry is broken. The curve labeled “0LL” shows the interaction energy obtained by projection to the LLL (no Landau-level mixing) for comparison [30]. Inset: schematic energetic competition between CDW and FCI states at $\nu_{h}=2/3$ and $3/5$, analogous to Landau-level systems [18, 21]. adjust magic angle (might be closer to $3.6$?)
@@
The coefficient of $B$ in this relation is sometimes referred to as the Středa slope [47]. As noted elsewhere [56, 14], the magnetic-field dependence of $p^{*}$ extracted experimentally does not follow a fixed-commensuration form like the equation above, and instead exhibits an apparent Středa slope intermediate between those of the competing FCI states at bracketing filling factors. This behavior is expected if the system responds to changes in magnetic field by changing the crystal unit-cell area $A^{\rm EC}$ (or by switching between nearby commensurations), and it requires that Wigner-crystal energies are less sensitive to $\nu_{h}$ than the competing FCI states, as indicated schematically in the caption of Fig. 2. As $B$ is varied, a sequence of reentrant integer quantum Hall Wigner-crystal states with differing $A^{\rm EC}$ can then be stabilized over the interval of $\nu_{h}$ between the $\nu_{h}=2/3$ and $\nu_{h}=3/5$ FCI Středa lines, where CDW states compete most effectively.
@@ I.1 Continuum model
We follow Refs. [46, 34, 23] closely for this section. Homobilyer TMDs can be written in a two layer effective model written in the layer basis [54]:
@@
Here, $\hat{\bm{k}}=-i\nabla$, is the wavenumber operator and $\bm{k}_{b/t}$ are the bottom and top momentum shifts corresponding to the top and bottom TMD layer $K$ (or $K^{\prime}$) points. $\Delta_{b/t}(\bm{r})$ are the intra-layer potentials and $\Delta_{T}(\bm{r})$ is the inter-layer tunneling term. We do a layer dependent gauge transformation:
$H_{TMD}\rightarrow\begin{pmatrix}e^{-i\bm{k}_{b}\cdot r}&0\\ 0&e^{-i\bm{k}_{t}\cdot r}\end{pmatrix}H_{TMD}\begin{pmatrix}e^{i\bm{k}_{b}\cdot r}&0\\ 0&e^{i\bm{k}_{t}\cdot r}\end{pmatrix}$. (S2)
This makes the resulting Hamiltonian quasi-periodic in the Moire unit cell (periodic up to a unitary transformation, Note that the spectrum remains periodic). Under this transformation:
@@
We focus on the first shell approximation for the intralayer potentials and interlayer tunelling:
@@
Where, $\bm{G}_{\bm{j}}=C_{6}^{j}\bm{G}_{1}$ for $j=1\dots 6$ and $\bm{G}_{1}=k_{\theta}(1/2,\sqrt{3}/2)$. $\bm{q}_{i+1}=C_{3}^{i}\bm{q}_{1}$,$i=0,1,2$ and use twist angle parameters obtained from large scale DFT calculations that were done for the twist angle $3.89^{\circ}$.
@@ I.2 Adiabatic model
We perform a local unitary transformation at each $\bm{r}$ such that diagonalizes the pseudospin texture term in Eq. (S4):
@@
$H^{\prime}=\begin{pmatrix}-\dfrac{1}{2m^{*}}\left(\hbar\hat{\bm{k}}+e\bm{A}_{\uparrow\uparrow}\right)^{2}-\dfrac{e^{2}}{2m^{*}}|\bm{A}_{\uparrow\downarrow}|^{2}+|\Delta(\bm{r})|+\Delta_{0}(\bm{r})&-\frac{\hbar e}{2m^{*}}\{\hat{\bm{k}},\bm{A}_{\uparrow\downarrow}\}-\frac{e^{2}}{2m^{*}}\bm{A}_{\uparrow\downarrow}(\bm{A}_{\uparrow\uparrow}+\bm{A}_{\downarrow\downarrow})\\ -\frac{\hbar e}{2m^{*}}\{\hat{\bm{k}},\bm{A}_{\downarrow\uparrow}\}-\frac{e^{2}}{2m^{*}}\bm{A}_{\downarrow\uparrow}(\bm{A}_{\uparrow\uparrow}+\bm{A}_{\downarrow\downarrow})&-\dfrac{1}{2m^{*}}\left(\hbar\hat{\bm{k}}+e\bm{A}_{\downarrow\downarrow}\right)^{2}-\dfrac{e^{2}}{2m^{*}}|\bm{A}_{\uparrow\downarrow}|^{2}-|\Delta(\bm{r})|+\Delta_{0}(\bm{r})\end{pmatrix}$. (S15)
We now project the Hamiltonian to the low-energy sector denoted by $\uparrow$. This is valid within the assumption the kinetic energy fluctuations are small $\hbar^{2}/2m^{*}A_{M}\ll|\Delta(\bm{r})|$. We then obtain the Adiabatic Hamiltonian:
@@
$U_{\mathrm{ad}}(\bm{r})$ $=-\dfrac{e^{2}}{2m^{*}}|A_{\uparrow\downarrow}|^{2}+|\Delta(\bm{r})|+\Delta_{0}(\bm{r}),$ (S18)
$A_{\uparrow\uparrow}$ $=\bm{\mathcal{A}}_{0}+\delta\bm{\mathcal{A}},\quad\nabla\times\bm{\mathcal{A}}_{0}=\mathcal{B}_{0}\hat{\bm{z}},\quad\nabla\times\delta\bm{\mathcal{A}}=\delta\mathcal{B}(\bm{r})\hat{\bm{z}}$. (S19)
@@
The scalar potential $U_{\mathrm{ad}}(\bm{r})$ collects contributions from $|\bm{\Delta}(\bm{r})|$, $\Delta_{0}(\bm{r})$, and the quantum metric of the pseudospin texture. The pseudo magnetic field is broken down into a spatially constant part $\mathcal{B}_{0}$ with flux quantum per moire unit cell and $\delta\mathcal{B}$ a spatially periodic (pseudo) magnetic field with zero spatial average within the Moire Unit cell. Note that Eq. (2) of the main text is written in Gaussian units. approximation.

### 10 Mar 2026 v2

@@ Abstract
The recent observations of fractional quantum anomalous Hall effects (FQAHEs) in twisted MoTe2 (tMoTe2) homobilayers [6, 40, 55, 59, 22] and in multilayer rhombohedral graphene [29, 36, 28] aligned with hexagonal boron nitride, have established these systems as platforms for fractional Chern insulator (FCI) states. FCIs emerge when strong electron–electron interactions dominate the kinetic energy within a topologically nontrivial, nearly flat electronic band [35, 45, 52, 44, 20, 4, 19]. Like their magnetic-field-driven counterparts — fractional quantum Hall effect (FQHE) [49] states — FCIs compete closely with interaction-driven phases that break translational symmetry [42, 24]. In the FQHE case, electron and hole Wigner crystals [53], bubble phases [16], and stripe phases [27] are well-established as ground states over some ranges of Landau-level filling [57, 58, 31, 9, 5, 30, 10]. The recent observation of reentrant integer quantum Hall states [56, 29, 38] in both rhombohedral graphene and TMD homobilayer systems suggests that broken translational symmetry states are also ground states at some fractional band filling factors in these systems.
@@
Figure 1: Flat bands, real-space structure and effective moiré potential of tMoTe2. (a) Bandwidth (meV) versus twist angle of the Adiabatic model (first-shell approximation) and the Continuum model for tMoTe2. (b) Schematic of the moiré unit cell illustrating MX, XM and MM sites, and $\sqrt{3}\times\sqrt{3}$ and $2\times 2$ commensurate supercells. (c) Amplitude of the leading (first-shell) harmonic of the effective moiré potential $V_{1}(\theta)$ versus twist angle; $V_{1}$ changes sign at $\theta_{c}\simeq 3.7^{\circ}$, switching the potential minima from MX (or XM, respectively) on the unit cell edges to $M$ at the unit-cell center. (d) The Landau-level mixing parameter, $\kappa\equiv[e^{2}/\varepsilon\,\ell_{B}]/(\hbar\omega_{c})$, versus twist angle for $\varepsilon=10.0$. In (a,d), the vertical gray line marks $\theta_{c}\simeq 3.7^{\circ}$. (e) Illustration of how the valence-band holes tend to localize on the MX or XM sites for $\theta<\theta_{c}$ and on the MM sites for $\theta>\theta_{c}$ in tMoTe2. We use model parameters from Ref. [51].
@@
where $\hat{\bm{\Pi}}=-\mathrm{i}\hbar\nabla+e\bm{\mathcal{A}}_{0}/c$. The scalar potential $U_{\mathrm{ad}}(\bm{r})$ collects contributions from $|\bm{\Delta}(\bm{r})|$, $\Delta_{0}(\bm{r})$, and the quantum metric of the pseudospin texture. The uniform component $\mathcal{B}_{0}$ is conveniently treated by using the corresponding Landau-level representation for $H_{\mathrm{ad}}$. The field $\mathcal{B}_{0}$ defines an effective magnetic length $\ell_{B}=\sqrt{\hbar c/(e\,\mathcal{B}_{0})}$ and cyclotron energy $\hbar\omega_{c}=e\,\mathcal{B}_{0}/(m^{*}c)\propto\theta^{2}$ . We quantify the interaction strength by the Landau-level mixing parameter $\kappa\equiv[e^{2}/(\varepsilon\,\ell_{B})]/(\hbar\omega_{c})\propto 1/\theta$. When projected to the lowest Landau-level (LLL), $V$ acts as an effective scalar potential $V(\bm{r})$, see Supplemental Material [3] for details. We denote the amplitude of its leading (first-shell) Fourier harmonic by $V_{1}(\theta)$, see Fig. 1(c). The sign change of $V_{1}$ at the critical angle $\theta_{c}$ reverses the landscape of potential minima and therefore selects the preferred CDW pinning site.
Interactions and Hartree–Fock theory.— We study the influence of long-range Coulomb interactions, $H_{\mathrm{C}}=\frac{1}{2A}\sum_{\bm{q}}V_{q}:\hat{n}_{-\bm{q}}\hat{n}_{\bm{q}}:$, using self-consistent Hartree–Fock (HF) theory, which provides an accurate description of CDW states at fractional band fillings. We employ the Landau gauge $\bm{\mathcal{A}}_{0}=\mathcal{B}_{0}\,x\,\hat{\bm{y}}$, in which states within a Landau-level are labeled by guiding center $X$. The mean-field Hamiltonian is
@@
CDW properties.— When the moiré potential is ignored, our problem reduces to that of holes in Landau levels for which the CDW states are hole Wigner crystals (one hole per unit cell) when $\nu_{h}<1/2$ and electron Wigner crystals for $\nu_{h}>1/2$  [^1] We work in a Hilbert space in which the microscopic degrees of freedom are holes. By electron Wigner crystal we mean a state in which the CDW unit cell area contains one hole less than a full Landau level. When Landau-level mixing is weak, these states are expected to be stable for $\nu_{h}>1/2$ because electrons are then more readily localized than holes.. The HF ground-state energies $E_{\mathrm{tot}}(\theta)$ we obtain for $\nu_{h}=1/3,\,2/3,\,1/4,\,1/2,$ and $3/4$ demonstrate that the moiré potential influences the CDW ground state. We first discuss $\nu_{h}=1/3$ and $2/3$, for which we consider $\sqrt{3}\!\times\!\sqrt{3}$ CDW supercells containing, respectively, one and two holes.
@@
Figure 2: Competing charge-ordered states over a range of hole fillings and twist angles. (a) Total energy per hole $E_{\textrm{tot}}$ versus twist angle $\theta$ for competing phases at fillings $\nu_{h}=1/4,\,1/3,\,1/2,\,2/3,\,3/4$. The line style indicates the rotational symmetry of the corresponding state ($C_{6}$, $C_{3}$, $C_{2}$); first-order transition angles are marked by square markers. Numbered markers label configurations whose real-space densities are shown in (b). (b) Real-space hole density $\rho_{h}(\bm{r})$ for the six labeled configurations in (a). White hexagons mark the moiré unit cell and red hexagons mark the CDW unit cell. Single-particle energies in (a) are relative to the twist-angle-dependent moiré band energy at $\nu_{h}=0$.
At $\nu_{h}=2/3$ the trend reverses: two holes fill each $\sqrt{3}\!\times\!\sqrt{3}$ CDW unit cell as illustrated in Fig. 2(a). For $\theta<\theta_{c}$, holes occupy both MX and XM sites and form a $C_{6}$-symmetric state, leaving MM sites to be occupied by electrons (missing holes). (see Fig. 2(b), panel 4.) For $\theta>\theta_{c}$, the moiré potential peaks at MX and XM sites and drives hole accumulation at MM, and either an MX or an XM site, lowering the symmetry from $C_{6}\!\to\!C_{3}$, corresponding to electrons at XM or MX sites (see Fig. 2(b), panel 5.) [1]. Note that the energy difference between these crystal motifs is much smaller at $\nu_{h}=2/3$ compared to at $\nu_{h}=1/3$. We attribute this difference to a reduction in the lattice-pinning energy in the electron crystal case that cannot take advantage of Landau-level mixing  [^2] At $\nu_{h}=1/3$ the crystal is a hole Wigner crystal in which the hole density $\rho_{h}(\bm{r})$ is maximized at lattice sites and increases with the Landau-level mixing parameter $\kappa$ (see Fig. 1(d)). In contrast, at $\nu_{h}=2/3$ the competing state is an electron CDW, corresponding to fully depleting $\rho_{h}(\bm{r})$ at lattice sites. Because $\rho_{h}(\bm{r})\!\geq\!0$, the hole density at lattice sites is limited and LL mixing is less effective in further lowering energies..
@@
Figure 3: (a) Lowest-Landau-level weight $w_{\mathrm{LLL}}$ of the self-consistent CDW solutions versus twist angle $\theta$ (see SM [3] for the definition of $w_{\mathrm{LLL}}$). The light-blue band shows the range across fillings $\nu_{h}\in\{1/4,1/3,1/2,2/3,3/4\}$ as a function of $\theta$. Representative curves are overlaid: color encodes filling $\nu_{h}$, and line style encodes symmetry ($C_{6}$, $C_{3}$, $C_{2}$). The light-gray curve traces the minimum $w_{\mathrm{LLL}}^{\mathrm{min}}(\theta)$ across all datasets. (b) Total energy per hole versus filling at representative twist angles $\theta=3.0^{\circ}$, $3.7^{\circ}$, and $4.2^{\circ}$; shaded bands cover the energy range over $\theta\in[2.5^{\circ},4.5^{\circ}]$. Marker style indicates when $C_{6}$ symmetry is broken. The curve labeled “0LL” shows the interaction energy obtained by projection to the LLL (no Landau-level mixing) for comparison [30]. Inset: schematic energetic competition between CDW and FCI states at $\nu_{h}=2/3$ and $3/5$, analogous to Landau-level systems [18, 21].
@@
The coefficient of $B$ in this relation is sometimes referred to as the Středa slope [47]. As noted elsewhere [56, 14], the magnetic-field dependence of $p^{*}$ extracted experimentally does not follow a fixed-commensuration form like the equation above, and instead exhibits an apparent Středa slope intermediate between those of the competing FCI states at bracketing filling factors. This behavior is expected if the system responds to changes in magnetic field by changing the crystal unit-cell area $A^{\rm EC}$ (or by switching between nearby commensurations), and it requires that Wigner-crystal energies are less sensitive to $\nu_{h}$ than the competing FCI states, as indicated schematically in the inset of Fig. 3(b). As $B$ is varied, a sequence of reentrant integer quantum Hall Wigner-crystal states with differing $A^{\rm EC}$ can then be stabilized over the interval of $\nu_{h}$ between the $\nu_{h}=2/3$ and $\nu_{h}=3/5$ FCI Středa lines, where CDW states compete most effectively.
@@ I.1 Continuum model
We follow Refs. [46, 34, 23] closely for this section. Homobilayer TMDs can be described by a two-layer effective model written in the layer basis [54]:
@@
Here, $\hat{\bm{k}}=-i\nabla$ is the wavenumber operator and $\bm{k}_{b/t}$ are the bottom and top momentum shifts corresponding to the bottom and top TMD layer $K$ (or $K^{\prime}$) points. $\Delta_{b/t}(\bm{r})$ are the intra-layer potentials and $\Delta_{T}(\bm{r})$ is the inter-layer tunneling term. We do a layer-dependent gauge transformation:
$H_{TMD}\rightarrow\begin{pmatrix}e^{-i\bm{k}_{b}\cdot\bm{r}}&0\\ 0&e^{-i\bm{k}_{t}\cdot\bm{r}}\end{pmatrix}H_{TMD}\begin{pmatrix}e^{i\bm{k}_{b}\cdot\bm{r}}&0\\ 0&e^{i\bm{k}_{t}\cdot\bm{r}}\end{pmatrix}$. (S2)
This makes the resulting Hamiltonian quasi-periodic in the moiré unit cell (periodic up to a unitary transformation, note that the spectrum remains periodic). Under this transformation:
@@
We focus on the first shell approximation for the intralayer potentials and interlayer tunneling:
@@
Where, $\bm{G}_{j}=C_{6}^{j}\bm{G}_{1}$ for $j=1\dots 6$ and $\bm{G}_{1}=k_{\theta}(1/2,\sqrt{3}/2)$. We define $\bm{q}_{i+1}=C_{3}^{i}\bm{q}_{1}$, $i=0,1,2$ and use twist angle parameters obtained from large-scale DFT calculations that were done for the twist angle $3.89^{\circ}$.
@@ I.2 Adiabatic model
We perform a local unitary transformation at each $\bm{r}$ such that it diagonalizes the pseudospin texture term in Eq. (S4):
@@
$H^{\prime}=\begin{pmatrix}-\dfrac{1}{2m^{*}}\left(\hbar\hat{\bm{k}}+e\bm{A}_{\uparrow\uparrow}\right)^{2}-\dfrac{e^{2}}{2m^{*}}|\bm{A}_{\uparrow\downarrow}|^{2}+|\bm{\Delta}(\bm{r})|+\Delta_{0}(\bm{r})&-\frac{\hbar e}{2m^{*}}\{\hat{\bm{k}},\bm{A}_{\uparrow\downarrow}\}-\frac{e^{2}}{2m^{*}}\bm{A}_{\uparrow\downarrow}(\bm{A}_{\uparrow\uparrow}+\bm{A}_{\downarrow\downarrow})\\ -\frac{\hbar e}{2m^{*}}\{\hat{\bm{k}},\bm{A}_{\downarrow\uparrow}\}-\frac{e^{2}}{2m^{*}}\bm{A}_{\downarrow\uparrow}(\bm{A}_{\uparrow\uparrow}+\bm{A}_{\downarrow\downarrow})&-\dfrac{1}{2m^{*}}\left(\hbar\hat{\bm{k}}+e\bm{A}_{\downarrow\downarrow}\right)^{2}-\dfrac{e^{2}}{2m^{*}}|\bm{A}_{\uparrow\downarrow}|^{2}-|\bm{\Delta}(\bm{r})|+\Delta_{0}(\bm{r})\end{pmatrix}$. (S15)
We now project the Hamiltonian to the low-energy sector denoted by $\uparrow$. This is valid under the assumption that the kinetic energy fluctuations are small $\hbar^{2}/2m^{*}A_{M}\ll|\bm{\Delta}(\bm{r})|$. We then obtain the Adiabatic Hamiltonian:
@@
$U_{\mathrm{ad}}(\bm{r})$ $=-\dfrac{e^{2}}{2m^{*}}|A_{\uparrow\downarrow}|^{2}+|\bm{\Delta}(\bm{r})|+\Delta_{0}(\bm{r}),$ (S18)
$\bm{A}_{\uparrow\uparrow}$ $=\bm{\mathcal{A}}_{0}+\delta\bm{\mathcal{A}},\quad\nabla\times\bm{\mathcal{A}}_{0}=\mathcal{B}_{0}\hat{\bm{z}},\quad\nabla\times\delta\bm{\mathcal{A}}=\delta\mathcal{B}(\bm{r})\hat{\bm{z}}$. (S19)
@@
The scalar potential $U_{\mathrm{ad}}(\bm{r})$ collects contributions from $|\bm{\Delta}(\bm{r})|$, $\Delta_{0}(\bm{r})$, and the quantum metric of the pseudospin texture. The pseudomagnetic field is decomposed into a spatially constant part $\mathcal{B}_{0}$ with one flux quantum per moiré unit cell and into $\delta\mathcal{B}$, a spatially periodic (pseudo)magnetic field with zero spatial average within the moiré unit cell. Note that Eq. (2) of the main text is written in Gaussian units.